# Generative AI for Low-Carbon Artificial Intelligence of Things

### v1

@@ IV-C3 Numerical results
We perform experiments on NVIDIA GeForce RTX 3080 Laptop GPU. The numerical result module (a) of Fig. 3 shows test reward curves of the proposed GDM-based algorithm and Proximal Policy Optimization (PPO) for optimal strategy design. We can observe that the GDM achieves higher test rewards than PPO, indicating better performance. The reason is that GDMs generate optimal strategies by diffusion process that can mitigate the impacts of noise and randomness[14]. The numerical result module (b) of Fig. 3 illustrates optimal strategies and the corresponding carbon emissions under different network environments. We can observe that due to the exploration experience during the denoising, the GDM-based algorithm can determine the strategy for low carbon emissions.
@@ VI Conclusion
In this article, we presented the prospect of GAI for low-carbon AIoT. First, we investigated the carbon challenges of mobile networks and systematically reviewed GAI techniques. Then, we explored the potential applications of GAI in enabling low-carbon AIoT. Inspired by the outstanding capabilities of LLMs, we proposed an LLM-enabled carbon emission optimization framework enabled by RAG, thus generating more accurate and reliable carbon emission optimization problems. Furthermore, we utilized GDMs to generate optimal strategies for carbon emission reduction. To validate the effectiveness of the proposed framework, we conducted a case study on mobile AIGC task offloading in a metaverse environment. Finally, we discussed potential research directions that can further achieve low-carbon AIoT.

### v2

@@ IV-C3 Numerical results
We perform experiments by using PyTorch on NVIDIA GeForce RTX 3080 Laptop GPU. The numerical result module (a) of Fig. 3 shows test reward curves of the proposed GDM-based algorithm and Proximal Policy Optimization (PPO) for optimal strategy design. We can observe that the GDM achieves higher test rewards than the PPO, indicating better performance. The reason is that GDMs generate optimal strategies by diffusion process that can mitigate the impacts of noise and randomness[5]. To demonstrate that training GDMs does not result in excessive carbon emissions, we use a Python package called CodeCarbon [^7] https://github.com/mlco2/codecarbon and estimate the power consumption and carbon emissions caused by GDM training to be $8.1488.148$ Wh and $1.6721.672$ g, respectively. The numerical result module (b) of Fig. 3 illustrates optimal strategies and the corresponding carbon emissions under different network environments. We can observe that due to the exploration experience during the denoising, GDMs can determine the optimal strategy for low carbon emissions.
@@ VI Conclusion
[…]